OpenStudy (anonymous):

x=9 y=3/5 write a direct variation equation

OpenStudy (anonymous):

y = kx where k is a constant to find value of k plug in values of x and y: 3/5 = k * 9 k = 3/5 / 9 = 3/45 = 1/15 equation is y = (1/15) x

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

nope thanks man